A home’s backyard is 20 feet by 15 feet. What are its perimeter and area?
Answer by jim_thompson5910(35256)

Area

A = L*W

A = 20*15

A = 300 square feet

=======================================================

Perimeter

P = 2(L + W)

P = 2*(20+15)

P = 2*35

P = 70 feet